1.Mr. Q is a 56 y.o male admitted to the hospital after a motor vehicle accident. His admission weight is 280 Ibs. what is his admission weight in kg?

2. Ms. F is being converted from an IV antibiotic to PO cefadroxil in preparation for her discharge home. order: cefadroxil 0.5 g PO bid. supply: cefadroxil 500 mg tablets. how many tablets will Ms. F be taking in a single dose?

3. The has an order for lorazepam (Ativan) 0.5 mg IV q6h PRN for agitation. Available is a vial containing 2 mg/ml. the nurse should draw up------- ml of solution for injection.

4. A client has a temperature of 101.2 F. There is an order for acetaminophen (Tylenol) 650 mg PO for fever, and 325 mg tabs are available. the nurse should give----- tablets
Answer #1

ANSWER.

1) Calculation of admission weight in kg.

       Weight of the patient in pounds = 280 lbs.

                    1 lbs = 0.454 kg.

      So

        Weight of the patient in kg = 280 * 0.454

                                                  = 127.1 kg.

2) Number of tablets Mr.F take in a single dose.

            Ordered dose = 0.5 gm bid

                                   = 500 mg (0.5 gm) bid (twice a day).

           Available dose = 500 mg /tab.

     Number of tablets required in a single dose = Ordered dose/Available dose

                                                                   = 500 / 500

                                                                   = 1 tablet (single dose).

3) Calculation of volume of solution for injection.

              Formula = Desired dose/Dose in hand * Quantity.

        Here

             Desired dose(D) = 0.5 mg.

             Dose in hand (H) = 2 mg.

                   Quantity (Q) = 1 ml.

As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 0.5/2 * 1

                                                        = 0.25 ml.

4) Number of tablets required.

            Ordered dose = 650 mg.

            Available dose = 325 mg/tab.

    Number of tablets required = Ordered dose/Available dose.

                                                = 650/325

                                                = 2 tablets.